Solve for r
a) A = P(1+r)^2
Can you please help me out? Thanks so much in advance:)
Can you also please show me the steps it would help me understand better:)

Answer by Cromlix(4381)

A = P(1 + r)^2
First of all square root both sides
this gets rid of the square on the right.
Square rootA = P(1 - r)
Divide both sides by P
This gets rid of the P
on the right
Square root A/ P = 1 - r
Take away 1 from both sides
This gets rid of the 1 on the right.
(square root A)/P - 1 = - r
Finally multiply both sides by -1
This makes r positive.
r = 1 - (square rootA)/ P
Hope this helps
:-)